# Probabilistic Emulation of a Global Climate Model with Spherical DYffusion

### v1

@@ Abstract
Data-driven deep learning models are on the verge of transforming global weather forecasting. It is an open question if this success can extend to climate modeling, where long inference rollouts and data complexity pose significant challenges. Here, we present the first conditional generative model able to produce global climate ensemble simulations that are accurate and physically consistent. Our model runs at 6-hourly time steps and is shown to be stable for 10-year-long simulations. Our approach beats relevant baselines and nearly reaches a gold standard for successful climate model emulation. We discuss the key design choices behind our dynamics-informed diffusion model-based approach which enables this significant step towards efficient, data-driven climate simulations that can help us better understand the Earth and adapt to a changing climate.
@@ 1 Introduction
Training relatively cheap-to-run data-driven surrogates to emulate global climate models could provide a compelling alternative. Although recent deep learning models are on the verge of transforming the conceptually similar field of medium-range weather forecasting [4, 31, 8, 41], these advances do not directly transfer to long-term climate projections. Indeed, most such models only report forecasts up to two weeks into the future and may diverge or become physically inconsistent over longer simulations. In contrast, climate projections demand accurate and stable simulations of the global Earth system spanning decades or centuries, requiring reliable reproduction of long-term statistics.
In Figure 1 we quantitatively show this divergence between the medium-range weather forecasting skill of ML models (measured as the average RMSE on 5-day forecasts) and their performance on longer climate time scales (measured as the RMSE of the 10-year time-mean). We have verified that this finding holds regardless of the analyzed variable and the proxy used for weather performance, which we discuss in more detail in Section D.4. Heuristically, optimizing weather skill ensures that a climate model takes a locally accurate path around the climate ’attractor’, but it does not guarantee that small but systematic errors may not build up to distort that simulated attractor to have biased long-term climate statistics. While this is a little-discussed observation in the ML community, the climate modeling community has documented it for physics-based models [13, 45].
@@
Our generative model is a leap forward toward purely ML-based climate projections that are both efficient and accurate. Our main contributions are:
@@
    We show that weather forecasting performance is not a strong indicator of long-term climate biases.
@@ 2 Related Work
However, advances in weather forecasting hardly transfer to long-term climate projections. Fully data-driven models fail to maintain stability beyond two-week-ahead forecasts, as errors accumulate over their autoregressive rollouts. Weyn et al. [59] and Bonev et al. [5] showed stable forecasts for horizons of up to six weeks and one year, respectively. Only recently, Watt-Meyer et al. [58] notably achieved stable and accurate 10-year simulations, followed by another deterministic SFNO-based climate emulator showing promising results using four prognostic variables [17]. Easier, but less flexible and informative, alternatives to full-scale temporal modeling of atmospheric dynamics, include emulation of annual means given an emission scenario [57, 36] or temporal super-resolution of monthly means [3].
@@
Diffusion models [20, 50, 51, 52] have demonstrated significant success in generating data such as natural images and videos. Their direct applications to autoregressive forecasting [29, 41] and downscaling [55, 34] of physical data have shown promising results. However, these approaches inherit the computational complexity associated with training and sampling from standard diffusion models. This is particularly prohibitive for autoregressive predictions on climate time scales, as the total number of neural network forward passes increases proportionally with the number of sampling steps, typically ranging from 20 to 1000. Consequently, recent research that leverages insights from diffusion models to balance predictive performance and sampling speed appears more promising for assessing their viability in climate simulations [48, 33]. Diffusion models have been largely used in tandem with a U-Net [46, 11] backbone architecture but recent work reveals promising results for using a vision transformer architecture for natural image synthesis tasks [40, 24, 19]. Our work explores a different, neural operator-based, architecture for Earth data.
@@ 5.1 Dataset and Experimental Setup
To compare our proposed method against ACE [58], we use the same dataset, training and evaluation setup. The dataset consists of 11 distinct 10-year-long simulations from the state-of-the-art global atmospheric model FV3GFS [63], saved every 6 hours. The forcings consist of annually repeating climatological sea surface temperature (1982-2012 average) and incoming solar radiation. Greenhouse gas and aerosol concentrations are kept fixed. The data was regridded conservatively from the cubed-sphere geometry of FV3GFS to a $1^{\circ}$ Gaussian grid, and filtered with a spherical harmonic transform round-trip to remove artifacts in the high latitudes. We train on 100 years of simulated data from FV3GFS, and evaluate the models on how well they can emulate a distinct 10-year-long validation simulation (i.e. $H=14600=10\times 365\times 4)$. The 11 simulations form an initial-condition ensemble, where each simulation is independent of the other–after some discarded spinup time–due to the chaoticity of the atmosphere [26]. For more details, see Appendix B.
@@ 5.2 Baselines
All models were trained on A6000 GPUs using distributed training on 2 up to 8 GPUs, ensuring that the effective batch size (number of GPUs $\times$ number of gradient accumulation steps $\times$ data loader batch size) remains the same (see Table 5). For a fair inference runtime comparison measuring the wall clock time needed to simulate 10 years (i.e. one full validation rollout), we run all deep-learning baselines on one A100 GPU. We also include the runtime for the physics-based FV3GFS climate model which was run on 96 cores of AMD EPYC 7H12 processors. The deep learning methods are not only much faster, but also much more energy-efficient than FV3GFS.
For illustrative purposes, we also report the complexity of a standard autoregressive diffusion model [20, 29, 41] approach in terms of the number of neural function evaluations (NFEs) needed to predict $h$ time steps, totaling to $Nh$ where $N$ is the number of sampling steps required to reverse the diffusion process. $N$ usually ranges from between 20 to 1000. This makes the use of such an approach unattractive for climate emulation since the resulting inference runtime would not offer as significant speed-ups over the physics-based reference model.
@@ 5.3 Climate Biases
The most crucial quality of an ML-based climate model is its ability to reproduce the climatology of the emulated reference system, i.e. the long-term average (“time-mean”) of weather states. The time-mean of the validation simulation is defined as $\frac{1}{H}\sum_{t=1}^{H}\boldsymbol{x}_{t}$. The time-mean for each model is defined as $\frac{1}{H}\sum_{t=1}^{H}\hat{\boldsymbol{x}}_{t}$, where $\hat{\boldsymbol{x}}_{t}$ is the model’s prediction for time step $t$. These two quantities are then compared against each other using the bias, i.e. prediction - target, and root mean squared error (RMSE) as key metrics of interest for analyzing climate biases. For the probabilistic methods, i.e. ours, DYffusion, and ACE-STO, we generate simulation ensembles by sampling from the model multiple times using the same initial conditions. Unless specified otherwise, all ensemble results are based on $M=25$ ensemble members. We evaluate these ensembles by computing the RMSE on the ensemble-mean prediction $\frac{1}{MH}\sum_{m=1}^{M}\sum_{t=1}^{H}\hat{\boldsymbol{x}}_{t,m}$ as well as on member-wise time-means $\frac{1}{H}\sum_{t=1}^{H}\hat{\boldsymbol{x}}_{t,m}$, where $m$ indexes the ensemble member dimension of the predictions. Standard deviations are computed over the respective member-wise errors. The corresponding “optimal noise floor” for the ML-based emulators is estimated by comparing the validation simulation with the 10-member reference ensemble. All metrics are area-weighted according to the size of the grid cell.
@@
Our method can be effectively ensembled to further reduce climate biases to $29.28\%$ on average across all variables. The ensemble-mean prediction (light shade) consistently improves time-mean RMSEs across all fields compared to using a single member’s prediction (dark shade in Fig. 2). In contrast, reducing climate biases through ensemble-averaging is challenging for ACE and ineffective for DYffusion and ACE-STO. More detailed results across additional fields are presented in Figure 6.
In Section D.2.3 and Figure 10, we show that ensembling our method not only excels in absolute terms but that it also reduces errors more dramatically than the probabilistic baselines as a function of the ensemble size relative to single-member performance. That is, our method generates more effective ensemble climate simulations. Still, there is some gap compared to the reference ensemble for some variables such as total water path.
@@
Here, we verify that our time-mean ensemble passes this challenging test, measured using the intra-ensemble variability of time-mean averages of a few important climate statistics simulated by ensembles consisting of 25 samples of our model vs. the ten reference simulations. We measure this variability by computing the standard deviation across the ensemble of time-means for each latitude-longitude grid point and then average it over all grid points. In Table 2 we show that the resulting global mean variability of the ensemble of time-means of our method is within 10-20% of those of the reference simulations for all tabulated variables (and other predicted fields). In Section D.2.2 we show that the corresponding global maps of the time-mean variability reveal similar spatial patterns. That is, our method generates ensemble climate simulations with decadal variability consistent with the underlying climate model.
@@ 6 Conclusion
We introduced an efficient diffusion model-based approach, complemented by a spherical-aware neural architecture for fully data-driven ensemble climate simulations. Our model generates stable, 10-year-long probabilistic predictions, consisting of 14600 time steps, with lower climate biases than relevant deterministic and probabilistic baselines. Our method’s climate biases nearly reach a gold standard for successful climate model emulation, marking a significant leap towards efficient, data-driven climate simulations that can help society better understand the Earth and adapt to a changing climate.
@@
To achieve real-world impact, the dataset will need to be expanded so that ML emulators can be evaluated (and trained) on climate change scenarios/simulations. This will require using time-varying climate change forcings such as greenhouse gas and aerosol concentrations. Thanks to the use of a state-of-the-art atmospheric model, FV3GFS, such data can be generated. Additionally, we only considered emulating the atmosphere, but to achieve a full Earth System Model (ESM) we also need to emulate (or couple to a physics-based model of) other components such as ocean, land, sea-ice, etc. Additionally, we have ignored output-only variables such as precipitation. Training separate prediction heads for these variables or extending the DYffusion framework to support them is an interesting direction for future work. It is important to stress that while our method is more than 25×25$\times$25 × faster than the reference physics-based climate model, it is still slower than deterministic models such as ACE. Physics-based models typically generate ensembles by perturbing both their inputs as well as internal model parameters. In this work, we focus on the latter point of model uncertainty by emulating a climate model with a diffusion model-based approach. Examining initial condition uncertainty is complementary to this approach and is left for future work.
@@ Acknowledgements
This work was supported in part by the U.S. Army Research Office under Army-ECASE award W911NF-23-1-0231, the U.S. Department Of Energy, Office of Science, IARPA HAYSTAC Program, CDC-RFA-FT-23-0069, DARPA AIE FoundSci, NSF Grants #2205093, #2146343, and #2134274.
S.R.C. acknowledges generous support from a summer internship and subsequent collaboration with the Allen Institute for AI (AI2), which is primarily funded by the estate of Paul G. Allen. We thank Zihao Zhou, Gideon Dresdner, and Peter Eckmann for providing helpful feedback on our work.
@@ Appendix B Dataset
In the subsections below, we elaborate on the dataset and variables that we use, including background information on FV3GFS and how it was configured in order to generate the training and validation data. Any listed data preprocessing steps below, are also described in appendix A from [58]. The final training and validation data can be downloaded from Google Cloud Storage following the instructions of the ACE paper at https://github.com/ai2cm/ace. The data are licensed under Creative Commons Attribution 4.0 International.
@@ B.1 Input, output and forcing variables
Symbol Description Units Time
$T_{k}$ Air temperature K Snapshot
$q^{T}_{k}$ Specific total water (vapor + condensates) kg/kg Snapshot
$u_{k}$ Wind speed in eastward direction m/s Snapshot
$v_{k}$ Wind speed in northward direction m/s Snapshot
$T_{s}$ Skin temperature of land or sea-ice K Snapshot
$p_{s}$ Atmospheric pressure at surface Pa Snapshot
@@
The complete list of input, output, and forcing variables used in this work is given in Table 3. The only difference to the work from [58] is that we do not consider diagnostic (output only) variables. The forcings consist of annually repeating climatological sea surface temperature (1982-2012 average), $T_{s}$, and incoming solar radiation, DSWRFTOA. Prescribed sea surface temperatures are simply “overwritten” on the skin temperature predictions of the ML models over all open ocean locations (when rolling out the ML-based simulation). The other forcing or input-only variables are added as an additional channel dimension. For evaluation, we also consider the derived variable called total water path which is computed as $\mathrm{TWP}=\frac{1}{g}\sum_{k}q_{k}^{T}\,dp_{k}$, i.e. as a function of surface pressure and the profile of specific total water.

### v2

@@ Abstract
Data-driven deep learning models are transforming global weather forecasting. It is an open question if this success can extend to climate modeling, where the complexity of the data and long inference rollouts pose significant challenges. Here, we present the first conditional generative model that produces accurate and physically consistent global climate ensemble simulations by emulating a coarse version of the United States’ primary operational global forecast model, FV3GFS. Our model integrates the dynamics-informed diffusion framework (DYffusion) with the Spherical Fourier Neural Operator (SFNO) architecture, enabling stable 100-year simulations at 6-hourly timesteps while maintaining low computational overhead compared to single-step deterministic baselines. The model achieves near gold-standard performance for climate model emulation, outperforming existing approaches and demonstrating promising ensemble skill. This work represents a significant advance towards efficient, data-driven climate simulations that can enhance our understanding of the climate system and inform adaptation strategies. [^1] Code is available at https://github.com/Rose-STL-Lab/spherical-dyffusion
@@ 1 Introduction
Training relatively cheap-to-run data-driven surrogates to emulate global climate models could provide a compelling alternative [15]. Although recent deep learning models are on the verge of transforming the conceptually similar field of medium-range weather forecasting [5, 38, 11, 51], these advances do not directly transfer to long-term climate projections [37]. Indeed, most such models only report forecasts up to two weeks into the future and may diverge or become physically inconsistent over longer simulations. In contrast, climate projections demand accurate and stable simulations of the global Earth system spanning decades or centuries, requiring reliable reproduction of long-term statistics.
In Figure 1 we quantitatively show this divergence between the medium-range weather forecasting skill of ML models (measured as the average RMSE on 5-day forecasts) and their performance on longer climate time scales (measured as the RMSE of the 10-year time-mean). We have verified that this finding holds regardless of the analyzed variable and the proxy used for weather performance, which we discuss in more detail in Section E.3. Heuristically, optimizing weather skill ensures that a climate model takes a locally accurate path around the climate ’attractor’, but it does not guarantee that small but systematic errors may not build up to distort that simulated attractor to have biased long-term climate statistics. While this is a little-discussed observation in the ML community, the climate modeling community has documented it for physics-based models [17, 54].
@@
Our generative model is a leap forward toward purely ML-based large ensemble climate projections that are both efficient and accurate. Our main contributions are:
@@
    We show that short-term weather performance does not necessarily translate to accurate reproduction of long-term climate statistics.
@@ 2 Related Work
However, advances in weather forecasting hardly transfer to long-term climate projections. Fully data-driven models fail to maintain stability beyond two-week-ahead forecasts, as errors accumulate over their autoregressive rollouts. Weyn et al. [68] and Bonev et al. [8] showed stable forecasts for horizons of up to six weeks and one year, respectively. Only recently, Watt-Meyer et al. [67] notably achieved stable and accurate 10-year simulations, followed by another deterministic SFNO-based climate emulator showing promising results using four prognostic variables [22]. Easier, but less flexible and informative, alternatives to full-scale temporal modeling of atmospheric dynamics, include emulation of annual means given an emission scenario [66, 30, 46, 42], temporal super-resolution of monthly means [4], or debiasing climate model output [3, 7, 45].
@@
Diffusion models [25, 59, 60, 61] have demonstrated significant success in generating data such as natural images and videos. While traditionally formulated for finite-dimensional spaces, these models have been extended to function spaces [40]. Their direct applications to autoregressive forecasting [35, 51] and downscaling [64, 43, 20] of physical data have shown promising results. However, these approaches inherit the computational complexity associated with training and sampling from standard diffusion models. This is particularly prohibitive for autoregressive predictions on climate time scales, as the total number of neural network forward passes increases proportionally with the number of sampling steps, typically ranging from 20 to 1000. Consequently, recent research that leverages insights from diffusion models to balance predictive performance and sampling speed appears more promising for assessing their viability in climate simulations [57, 41]. While diffusion models traditionally rely on U-Net architectures [55, 13], vision transformers have shown promising results in image synthesis [50, 29, 24]. Our work explores a different, neural operator-based, architecture for Earth data.
@@ 5.1 Dataset and Experimental Setup
To compare our proposed method against ACE [67], we use the same dataset, training and evaluation setup. The dataset consists of 11 distinct 10-year-long simulations from the state-of-the-art global atmospheric model FV3GFS [73], saved every 6 hours. The forcings consist of annually repeating climatological sea surface temperature (1982-2012 average) and incoming solar radiation. Greenhouse gas and aerosol concentrations are kept fixed. The data was regridded conservatively from the cubed-sphere geometry of FV3GFS to a $1^{\circ}$ Gaussian (i.e. $180\times 360$) grid, and filtered with a spherical harmonic transform round-trip to remove artifacts in the high latitudes. We train on 100 years of simulated data from FV3GFS, and evaluate the models on how well they can emulate a distinct 10-year-long validation simulation (i.e. $H=14600=10\times 365\times 4)$. The 11 simulations form an initial-condition ensemble, where each simulation is independent of the other–after some discarded spinup time–due to the chaoticity of the atmosphere [32]. For more details, see Appendix B.
@@ 5.2 Baselines
All models were trained on A6000 GPUs using distributed training on 2 up to 8 GPUs, ensuring that the effective batch size remains the same (see Figure 8). For a fair inference runtime comparison measuring the wall clock time needed to simulate 10 years (i.e. one full validation rollout), we run all deep-learning baselines on one A100 GPU. We also include the runtime for the physics-based FV3GFS climate model which was run on 96 cores (24 cores for the 2×2$\times$2 × coarser version) of AMD EPYC 7H12 processors. The deep learning methods are not only much faster, but also much more energy-efficient than FV3GFS.
For illustrative purposes, we also report the complexity of a standard autoregressive diffusion model [25, 35, 51] approach in terms of the number of neural function evaluations (NFEs) needed to predict $h$ time steps, totaling to $Nh$ where $N$ is the number of sampling steps required to reverse the diffusion process. $N$ usually ranges from between 20 to 1000. This makes the use of such an approach less attractive for climate emulation since the resulting inference runtime would not offer as significant speed-ups over the physics-based reference model.
@@ 5.3 Climate Biases
The most crucial quality of an ML-based climate model is its ability to reproduce the climatology of the emulated reference system, i.e. the long-term average (“time-mean”) of weather states. The time-mean of the validation simulation is defined as $\frac{1}{H}\sum_{t=1}^{H}\boldsymbol{x}_{t}$. The time-mean for each model is defined as $\frac{1}{H}\sum_{t=1}^{H}\hat{\boldsymbol{x}}_{t}$, where $\hat{\boldsymbol{x}}_{t}$ is the model’s prediction for time step $t$. These two quantities are then compared against each other using the bias, i.e. prediction - target, and root mean squared error (RMSE) as key metrics of interest for analyzing climate biases. For the probabilistic methods, i.e. ours, DYffusion, and ACE-STO, we generate simulation ensembles by sampling from the model multiple times using the same initial conditions. Unless specified otherwise, all ensemble results are based on $E=25$ ensemble members. We evaluate the ensemble performance using two metrics: the RMSE of the ensemble-mean prediction ($\frac{1}{EH}\sum_{e=1}^{E}\sum_{t=1}^{H}\hat{\boldsymbol{x}}_{t,e}$) and the RMSE of member-wise time-means ($\frac{1}{H}\sum_{t=1}^{H}\hat{\boldsymbol{x}}_{t,e}$), where $e$ indexes individual ensemble members. For the latter, standard deviations are computed over the member-wise errors. The corresponding “optimal noise floor” for the ML-based emulators is estimated by comparing the validation simulation with the 10-member reference ensemble. All metrics, which are fully defined in Appendix D, are weighted by grid cell area. It is important to acknowledge the potential for improving the estimate of the “noise floor” based on statistical significance testing and improved metrics [21].
@@
Ensemble averaging significantly enhances our method’s performance, reducing climate biases by $29.28\%$ on average across all variables. As shown by the light shading in Fig. 2, the ensemble-mean predictions consistently achieve lower time-mean RMSEs compared to single-member predictions (dark shading). This ensemble-based improvement distinguishes our approach from ACE-STO and DYffusion, where ensemble averaging proves less effective, and from ACE, where initial-condition perturbations would be required for ensembling. Additional results for more fields are available in Figure 9 of the Appendix. Our comprehensive evaluation in Table 4 includes ensemble metrics such as the Continuous Ranked Probability Score (CRPS) and spread-skill ratio. The results demonstrate that our method outperforms alternatives in emulating the 10-year time-mean climatology of the reference model for most variables and metrics. However, some challenges remain, particularly in matching the reference ensemble’s performance for stratospheric (level 0) variables and in achieving better ensemble scores.
@@
In Section E.4, we visualize two sample 10-year trajectories simulated by Spherical DYffusion as well as the corresponding validation simulation from FV3GFS. Supplementary videos demonstrate the full temporal evolution of key derived variables: near-surface wind speed [^3] https://youtu.be/7lHra7gBiBo and total water path [^4] https://youtu.be/Hac_xGsJ1qY. The emulated fields demonstrate high realism, closely mimicking the patterns and variability observed in actual climate model outputs. This showcases Spherical DYffusion’s capability to generate plausible and physically consistent climate scenarios over decadal timescales.
@@
Here, we verify that our time-mean ensemble passes this challenging test, measured using the intra-ensemble variability of time-mean averages of a few important climate statistics simulated by 25-member ensembles of the emulators vs. the ten reference simulations. We measure this variability by computing the area-weighted average of the standard deviation of time-means across the ensemble dimension. In Table 2 we show that the resulting global mean variability of the ensemble of time-means of our method is within 10-20% of those of the reference simulations for all tabulated variables (and other predicted fields). DYffusion achieves similarly accurate ensemble variability, while ACE-STO In Section E.1.2 we show that the corresponding global maps of the time-mean variability reveal similar spatial patterns. That is, our method generates ensemble climate simulations with decadal variability consistent with the underlying climate model.
100-year-long simulation.
We evaluate the long-term stability of Spherical DYffusion through a 100-year simulation, a critical timescale for many climate modeling applications. Figure 6 demonstrates the model’s robustness through time series of key global mean variables from a single (random) simulation, which completed in approximately 26 hours of wall-clock time. The model generates physically consistent temporal patterns in response to annually repeating forcings. Notably, Spherical DYffusion exhibits improved variability patterns compared to the baseline ACE model, which suffers from unrealistic annual fluctuations (e.g. see surface pressure).
Figure 6: Comparison of 100-year global mean simulations between Spherical DYffusion and ACE. From top to bottom: near-surface air temperature ($T_{7}$), total water path (TWP), and surface pressure ($p_{s}$). Both models are driven by identical annually repeating forcings. Spherical DYffusion demonstrates more stable trajectories, particularly evident in the surface pressure predictions, while maintaining physically realistic variability patterns. The consistent behavior across all variables indicates the model’s robustness for long-term climate simulations.
@@ 6 Conclusion
We introduce Spherical DYffusion, a novel approach that combines efficient diffusion modeling with a spherical-aware neural architecture to probabilistically emulate complex global climate dynamics across decadal to centennial timescales. Our model achieves lower climate biases than relevant deterministic and probabilistic baselines, getting significantly closer to the optimal performance provided by the emulated climate model. For climate model emulation problems, our approach presents a unique solution for balancing generative modeling, computational efficiency, and low climate biases. This opens up the ability to perform fully data-driven ensemble climate simulations.
@@
To achieve real-world impact, the dataset will need to be expanded so that ML emulators can be evaluated (and trained) on climate change scenarios/simulations. This will require using time-varying climate change forcings such as greenhouse gas and aerosol concentrations. Although our use of the state-of-the-art FV3GFS atmospheric model enables generation of such training data, any emulator will inherently reflect biases present in the base model. Additionally, we only considered emulating the atmosphere, but to achieve a full Earth System Model (ESM) we also need to emulate (or couple to a physics-based model of) other components such as ocean, land, sea-ice, etc. It is important to stress that while our method is more than 25×25$\times$25 × faster than the reference physics-based climate model, it is still slower than deterministic emulators such as ACE. Though our method characterizes model uncertainty through its generative design, extending it to incorporate initial condition uncertainty—a key component of traditional ensemble physics-based models—could further enhance its capabilities. The method also needs extension to handle output-only variables like precipitation, either through dedicated prediction heads or modifications to the DYffusion framework.
@@ Acknowledgements
This work was supported in part by the U.S. Army Research Office under Army-ECASE award W911NF-23-1-0231, the U.S. Department Of Energy, Office of Science, IARPA HAYSTAC Program, CDC-RFA-FT-23-0069, DARPA AIE FoundSci, DARPA YFA, NSF Grants #2205093, #2100237,#2146343, and #2134274. S.R.C. acknowledges generous support from a summer internship and subsequent collaboration with the Allen Institute for AI (Ai2), which is primarily funded by the estate of Paul G. Allen. We are grateful to Zihao Zhou, Gideon Dresdner, and Peter Eckmann for their insightful feedback, and to the anonymous reviewers for their constructive comments and valuable suggestions that helped strengthen this work.
@@ Appendix B Dataset
In the subsections below, we elaborate on the dataset and variables that we use, including background information on FV3GFS and how it was configured in order to generate the training and validation data. Any listed data preprocessing steps below are also described in appendix A from [67]. The final training and validation data can be downloaded from Google Cloud Storage following the instructions of the ACE paper at https://zenodo.org/records/10791087. The data are licensed under Creative Commons Attribution 4.0 International.
@@ B.1 Input, output and forcing variables
Symbol Description Units Time Is 3D?
$T_{k}$ Air temperature K Snapshot Yes
$q^{T}_{k}$ Specific total water (vapor + condensates) kg/kg Snapshot Yes
$u_{k}$ Wind speed in eastward direction m/s Snapshot Yes
$v_{k}$ Wind speed in northward direction m/s Snapshot Yes
$T_{s}$ Skin temperature of land or sea-ice K Snapshot No
$p_{s}$ Atmospheric pressure at surface Pa Snapshot No
@@
Derived, evaluation-only, variables
Symbol Description Units Time Is 3D?
$\mathrm{WS}_{k}$ Wind speed m/s Snapshot Yes
$\mathrm{TWP}$ Total water path mm Snapshot No
The complete list of input, output, and forcing variables used in this work is given in Table 3. The only difference to the work from [67] is that we do not consider diagnostic (output only) variables. The forcings consist of annually repeating climatological sea surface temperature (1982-2012 average), $T_{s}$, and incoming solar radiation, DSWRFTOA. Prescribed sea surface temperatures are simply “overwritten” on the skin temperature predictions of the ML models over all open ocean locations (when rolling out the ML-based simulation). The other forcing or input-only variables are added as an additional channel dimension. Derived variables are computed from the (predicted) prognostic variables as described below.
Derived variables.
For evaluation, we also consider the derived variable called total water path which is computed as $\mathrm{TWP}=\frac{1}{g}\sum_{k}q_{k}^{T}\,dp_{k}$, i.e. as a function of surface pressure and the profile of specific total water. Its units are $mm$ (or $kg/m^{2}$, assuming that water has a density of 1000 $kg/m^{3}$). The derived wind speed variable for level $k$ is computed based on the simulated meridional and zonal wind variables as $\mathrm{WS}_{k}=\sqrt{u_{k}^{2}+v_{k}^{2}}$. Its units are $m/s$.